# Defect versus defect: stationary states of single file marching in periodic landscapes with road blocks

### v2

@@ Abstract
Since a DW due to PD can be inside $T_{2}$, whereas a DW due to ED cannot be, an intriguing situation can arise when an LDW due to PD is formed in $T_{2}$, with $J_{\text{PD}}<J_{\text{ED}}$ only slightly. By tuning $(p,q)$, one can make $J_{\text{PD}}>J_{\text{ED}}$ again only slightly, for which an LDW is formed due to ED, which however cannot be in $T_{2}$! Thus the LDW position can jump discontinuously under a small change in the defect strengths; see movies 5 and 6 in SM sm . As an example, consider $n=1/2$ and $p_{c},\,q_{c}$ satisfy $J_{\text{PD}}=J_{\text{ED}}$ together with $\epsilon_{1}=\epsilon_{2}=1/3$. Then if $p=p_{c}-\delta_{1},\,q=q_{c}+\delta_{2},\,\delta_{1},\,\delta_{2}>0$ (but $\delta_{1},\,\delta_{2}$ are very small) $J_{\text{PD}}<J_{\text{ED}}$, we have an LDW at $x_{w}=1/2$ due to the point defect sm . On the other hand, if $\delta_{1},\,\delta_{2}<0$, $J_{\text{ED}}<J_{\text{PD}}$ and hence an LDW at $x_{w}=0$. Thus, as one moves ${\cal O}(\delta),\,\delta=\delta_{1},\,\delta_{2}$ being vanishingly small, the DW position changes by an ${\cal O}(1)$ amount - from $x_{w}=0$ to $x_{w}=1/2$. Thus $x_{w}$ as a function of $\delta_{1},\delta_{2}$ (or $p,q$) shows a jump across the line $J_{\text{PD}}=J_{\text{LD}}$: let $\delta\equiv 2\sqrt{\delta_{1}^{2}+\delta_{2}^{2}}$ be the “distance” between the two points $(p=p_{c}-\delta_{1},\,q=q_{c}+\delta_{2})$ and $p=p_{c}+\delta_{1},\,q=q_{c}-\delta_{2}$ in the $p-q$ plane. Then, $dx_{w}/d\delta$ diverges for small $\delta$. Such divergences, a hallmark of defect competition, can occur only when the two points in the $p-q$ plane lie on the two sides of the line determined by the condition $J_{\text{PD}}=J_{\text{ED}}$. This jump in $x_{w}$ across the line $J_{\text{PD}}=J_{\text{ED}}$ indicates a discontinuous transition between the PD and ED dominated steady states each with one LDW, with $x_{w}$ appearing as an order parameter.
The density profiles in LD-LD-LD and HD-HD-HD phases are straightforward to obtain by using PNC and current conservation; see SM sm . Phase boundaries (4), (6), (8) and (9) give the phase diagram in Fig. (2).
@@
In summary, we have developed a theory for bottleneck competition in closed, driven SFM by studying a conceptual model consisting of a TASEP with a PD and an ED in a ring geometry. This reveals how a dominant defect enforces a particular form of LDW, when sufficient number of particles are available. When they are competing, a pair of DDWs are obtained, instead of a single LDW. We show how the competition between the defects manifests strikingly in the discontinuous jump of the LDW location, as the system passes from being controlled by one defect to another. Our theory can be generalised to arbitrary number of PD or ED straightforwardly. Our results can be experimentally studied in model experiments on the collective motion of driven particles with light-induced activity light through a closed narrow circular channel channel1 ; channel2 . Suppression of rotational diffusion, e.g., by choosing ellipsoidal particles with the channel width shorter than the long axis of the particle everywhere, or by using dimer particles can enforce unidirectional movements. Lastly, in spite of the simplicity of our model, the results from it should give insight about the effects of bottlenecks on directional motion in more complex living ped1 ; ped2 ; ped3 or in-vitro coll1 ; coll2 ; coll3 systems.
@@
Supplemental Material
Defect versus defect: stationary states of single file marching in periodic landscapes with road blocks
I Mean-field theory for the stationary density profiles
In this Section, we set up and analyze the MFT equations blythe-sm . Let $n_{i}$ be the occupation at site $i$. In the MFT approximation, correlations are neglected and averages of products are replaced by products of averages blythe . While MFT is an adhoc approximation, it provides a good analytically tractable guideline to the steady state densities and phase diagrams. It is convenient to consider the system to be composed of three segments - $T_{1}$ between $i=1$ to $\epsilon_{1}L$, $T_{2}$ between $\epsilon_{1}L$ to $(\epsilon_{1}+\epsilon_{2})L$ and $T_{3}$ between $(\epsilon_{1}+\epsilon_{2})L$ to $L$, connected serially. The MFT equation for the density $n_{i}$ in the different segments reads
@@ I.1 LD-LD-LD phase
So far we have not imposed any conditions of the LD-LD-LD phase on the solutions (21). The pertinent question then is: Which of the two solutions in (21) is to be considered as the LD-LD-LD phase density solution? To settle this, we use the fact that in the limiting case with vanishing particle number, i.e., with $n\rightarrow 0$, the LD-LD-LD phase solution must smoothly go to zero. This consideration allows us to pick the right solution in (21) for the LD-LD-LD phase: we choose the solution that vanishes as $n\rightarrow 0$. Which one among the two in (21) does that depends upon the signs of $A$ and $B$, i.e., will be decided by $\epsilon_{2}$ and $n$. We thus note that the point defect has no macroscopic effect on the steady state density profiles. Instead at $x=0$, the location of the point defect, there is a local peak of height $h$ with vanishing width in the thermodynamic limit, where $h=\rho_{1}(1-p)/p$, such that the local density at $x=0$ is $\rho_{1}+h$ erwin-defect-sm ; niladri-tasep-sm . This essentially acts as a boundary layer between $T_{1}$ and $T_{3}$. Steady state density profiles in the LD-LD-LD phase with $n=0.15,p=0.2,\,q=0.45$ and $n=0.15,p=0.15,\,q=0.6$ with a system size $L=9000$ and $\epsilon_{1}=\epsilon_{2}=1/3$ are shown in Fig. 5. Good agreement between MFT and MCS results are observed.
Refer to captionRefer to caption
Figure 5: LD-LD-LD density plots with $n=0.15$: (left) $p=0.2,\,q=0.45$, (right) $p=0.15,\,q=0.6$. MFT (blue lines) and MCS (red points) results are shown. Good agreement between MFT and MCS results are observed.
I.2 HD-HD-HD phase
@@
Similar to the LD-LD-LD phase, the point defect has no macroscopic effect on the stationary densities. Instead, one has a local dip of depth $h^{\prime}$, with vanishing width in the thermodynamic limit, where $h^{\prime}=(1-\rho_{1})(1-p)/p$, such that at $x=0$, the density is $\rho_{1}-h^{\prime}$. Steady state density profiles in the HD-HD-HD phase with $n=0.8,p=0.2,\,q=0.45$ and $n=0.8,p=0.15,\,q=0.6$ with a system size $L=9000$ and $\epsilon_{1}=\epsilon_{2}=1/3$ are shown in Fig. 6. Good agreement between MFT and MCS results are observed.
Refer to captionRefer to caption
Figure 6: HD-HD-HD density plots: (left) $n=0.8,\,p=0.2,\,q=0.45$, (right) $n=0.8,\,p=0.15,\,q=0.6$. MFT (blue lines) and MCS (red points) results are shown. Good agreement between MFT and MCS results are observed.
I.3 MC-MC-MC phase
@@
II LDWs due to the point defect
@@
When the LDW is inside $T_{2}$, $T_{1}$ is in its HD phase, hence $\rho_{1}=\rho_{\text{HD}}$, and $T_{3}$ is in its LD phase, with $\rho_{3}=\rho_{\text{LD}}$. Further, since the LDW is located inside $T_{2}$, we must have $\rho_{2}=\rho_{+}$ for $\epsilon_{1}<x<x_{w}$; $\rho_{2}=\rho_{-}$ for $x_{w}<x<\epsilon_{1}+\epsilon_{2}$. Applying PNC we get
@@
Next, consider an LDW in $T_{3}$. Thus, $\rho_{3}(x)=\rho_{\text{HD}},\;\epsilon_{1}+\epsilon_{2}<x<x_{w}$, $\rho_{3}(x)=\rho_{\text{LD}},\;x_{w}<x<1$. Further, $\rho_{1}(x)=\rho_{\text{HD}}$, $\rho_{2}(x)=\rho_{+}$, since both $T_{1}$ and $T_{2}$ are in their HD phases. Then applying PNC, we obtain
@@
Our MFT and MCS results on LDWs in $T_{1},\,T_{3}$ are shown in Fig. 7 (left) and Fig. 7 (right) respectively. We find reasonable agreement between our MFT and MCS results.
Refer to captionRefer to caption
Figure 7: Plot of an LDW in (left) $T_{1}$ in the point defect dominated regime with $p=0.2,n=0.35,q=0.65,L=9000,\epsilon_{1}=\epsilon_{2}=1/3$, and (right) $T_{3}$ in the point defect dominated regime with $p=0.2,n=0.65,q=0.65,L=9000,\epsilon_{1}=\epsilon_{2}=1/3$. MFT (blue lines) and MCS (red points) results are shown (see text).
III LDWs due to the extended defect
@@
$(\epsilon_{1}-x_{w})\rho_{-}+x_{w}\rho_{+}+\frac{\epsilon_{2}}{2}+(1-\epsilon_{1}-\epsilon_{2})\rho_{+}=n.$ (37)

### v3

@@ III Domain Walls
Since a DW due to PD can be inside $T_{2}$, whereas a DW due to ED cannot be, an intriguing situation can arise when an LDW due to PD is formed in $T_{2}$, with $J_{\text{PD}}<J_{\text{ED}}$ only slightly. By tuning $(p,q)$, one can make $J_{\text{PD}}>J_{\text{ED}}$ again only slightly, for which an LDW is formed due to ED, which however cannot be in $T_{2}$! Thus the LDW position can jump discontinuously under a small change in the defect strengths; see movies 5 and 6 in SM sm . As an example, consider $n=1/2$ and $p_{c},\,q_{c}$ satisfy $J_{\text{PD}}=J_{\text{ED}}$ together with $\epsilon_{1}=\epsilon_{2}=1/3$. Then if $p=p_{c}-\delta_{1},\,q=q_{c}+\delta_{2},\,\delta_{1},\,\delta_{2}>0$ (but $\delta_{1},\,\delta_{2}$ are very small) $J_{\text{PD}}<J_{\text{ED}}$, we have an LDW at $x_{w}=1/2$ due to the point defect. On the other hand, if $\delta_{1},\,\delta_{2}<0$, $J_{\text{ED}}<J_{\text{PD}}$ and hence an LDW at $x_{w}=0$. Thus, as one moves ${\cal O}(\delta),\,\delta=\delta_{1},\,\delta_{2}$ being vanishingly small, the DW position changes by an ${\cal O}(1)$ amount - from $x_{w}=1/2$ to $x_{w}=0$. Thus $x_{w}$ as a function of $\delta_{1},\delta_{2}$ (or $p,q$) shows a jump across the line $J_{\text{PD}}=J_{\text{LD}}$: let $\delta\equiv 2\sqrt{\delta_{1}^{2}+\delta_{2}^{2}}$ be the “distance” between the two points $(p=p_{c}-\delta_{1},\,q=q_{c}+\delta_{2})$ and $(p=p_{c}+\delta_{1},\,q=q_{c}-\delta_{2})$ in the $p-q$ plane. Then, $dx_{w}/d\delta$ diverges for small $\delta$. Such divergences, a hallmark of defect competition, can occur only when the two points in the $p-q$ plane lie on the two sides of the line determined by the condition $J_{\text{PD}}=J_{\text{ED}}$. This jump in $x_{w}$ across the line $J_{\text{PD}}=J_{\text{ED}}$ indicates a hitherto unknown discontinuous nonequilibrium transition between the PD and ED dominated steady states each with one LDW, with $x_{w}$ appearing as an order parameter.
Phase boundaries (4), (6), (8) and (9) give the phase diagram in Fig. (2) $\epsilon_{1}=\epsilon_{2}=1/3$. For other values of $\epsilon_{1},\epsilon_{2}$, the phase boundaries change mildly, keeping the topology of the phase diagram unchanged; see Appendix D for phase diagrams with (i) $\epsilon_{1}=2/5,\epsilon_{2}=1/5$ and (ii) $\epsilon_{1}=1/4,\epsilon_{2}=1/2$.
IV Summary and outlook
We have thus developed a theory for bottleneck competition in closed, driven single file motion by studying a conceptual model consisting of a periodic TASEP with a point and an extended defects. We show how a dominant defect enforces a particular form of LDW, when sufficient number of particles are available. For competing defects, a pair of DDWs are obtained, instead of a single LDW. The competition between the defects manifests strikingly in the discontinuous jump of the LDW location, as the system passes from being controlled by one defect to another. Our theory can be generalised to arbitrary number of point or extended defects straightforwardly. Our MCS studies reveal some finite-size effects and also quantitative disagreement between the MFT and MCS predictions for moderate densities in the point defect dominated phase space region [see Fig. 3 (left) and Fig. 4 (left)]. This can be systematically studied in future by using the finite-size scaling analysis developed in Ref. soh . In contrast, no such disagreements are found in the extended defect dominate phase space regions. These contrasting behaviours may be understood heuristically in terms of an ED working as a particle reservoir for the remaining TASEP, weakening the effects of number conservation; see Ref. tirtha-anjan in a partly related model; see also Ref. tirtha-anjan-2 for similar issues.
@@ IV Summary and outlook
Our results can be experimentally studied in model experiments on the collective motion of driven particles with light-induced activity light through a closed narrow circular channel channel1 ; channel2 . Suppression of rotational diffusion, e.g., by choosing ellipsoidal particles with the channel width shorter than the long axis of the particle everywhere, or by using dimer particles can enforce unidirectional movements. Lastly, in spite of the simplicity of our model, the results from it should give insight about the effects of bottlenecks on directional motion in more complex living ped1 ; ped2 ; ped3 or in-vitro coll1 ; coll2 ; coll3 systems.
@@
Appendix A Mean-field theory for the stationary density profiles
In this Section, we set up and analyze the MFT equations blythe and use them to obtain the LD-LD-LD and HD-HD-HD phase densities. Let $n_{i}$ be the occupation at site $i$. In the MFT approximation, correlations are neglected and averages of products are replaced by products of averages blythe . While MFT is an adhoc approximation, it provides a good analytically tractable guideline to the steady state densities and phase diagrams. It is convenient to consider the system to be composed of three segments - $T_{1}$ between $i=1$ to $\epsilon_{1}L$, $T_{2}$ between $\epsilon_{1}L$ to $(\epsilon_{1}+\epsilon_{2})L$ and $T_{3}$ between $(\epsilon_{1}+\epsilon_{2})L$ to $L$, connected serially. The MFT equation for the density $n_{i}$ in the different segments reads
@@ A.1 LD-LD-LD phase
So far we have not imposed any conditions of the LD-LD-LD phase on the solutions (21). The pertinent question then is: Which of the two solutions in (21) is to be considered as the LD-LD-LD phase density solution? To settle this, we use the fact that in the limiting case with vanishing particle number, i.e., with $n\rightarrow 0$, the LD-LD-LD phase solution must smoothly go to zero. This consideration allows us to pick the right solution in (21) for the LD-LD-LD phase: we choose the solution that vanishes as $n\rightarrow 0$. Which one among the two in (21) does that depends upon the signs of $A$ and $B$, i.e., will be decided by $\epsilon_{2}$ and $n$. We thus note that the point defect has no macroscopic effect on the steady state density profiles. Instead at $x=0$, the location of the point defect, there is a local peak of height $h$ with vanishing width in the thermodynamic limit, where $h=\rho_{1}(1-p)/p$, such that the local density at $x=0$ is $\rho_{1}+h$ niladri-tasep ; erwin-defect . This essentially acts as a boundary layer between $T_{1}$ and $T_{3}$. Steady state density profiles in the LD-LD-LD phase with $n=0.15,p=0.2,\,q=0.45$ and $n=0.15,p=0.15,\,q=0.6$ with a system size $L=9000$ and $\epsilon_{1}=\epsilon_{2}=1/3$ are shown in Fig. 6 (top) and Fig. 6 (bottom) respectively. Good agreement between MFT and MCS results are observed.
Figure 6: LD-LD-LD density plots with $n=0.15$: (top) $p=0.2,\,q=0.45$, (bottom) $p=0.15,\,q=0.6$. MFT (blue lines) and MCS (red points) results are shown. Good agreement between MFT and MCS results are observed.
A.2 HD-HD-HD phase
@@
Similar to the LD-LD-LD phase, the point defect has no macroscopic effect on the stationary densities. Instead, one has a local dip of depth $h^{\prime}$, with vanishing width in the thermodynamic limit, where $h^{\prime}=(1-\rho_{1})(1-p)/p$, such that at $x=0$, the density is $\rho_{1}-h^{\prime}$. Steady state density profiles in the HD-HD-HD phase with $n=0.8,p=0.2,\,q=0.45$ and $n=0.8,p=0.15,\,q=0.6$ with a system size $L=9000$ and $\epsilon_{1}=\epsilon_{2}=1/3$ are shown in Fig. 7 (top) and Fig. 7 (bottom) respectively. Good agreement between MFT and MCS results are observed.
Figure 7: HD-HD-HD density plots: (top) $n=0.8,\,p=0.2,\,q=0.45$, (bottom) $n=0.8,\,p=0.15,\,q=0.6$. MFT (blue lines) and MCS (red points) results are shown. Good agreement between MFT and MCS results are observed.
A.3 MC-MC-MC phase
@@
Appendix B LDWs due to the point defect
@@
When the LDW is inside $T_{2}$, $T_{1}$ is in its HD phase, hence $\rho_{1}=\rho_{\text{HD}}$, and $T_{3}$ is in its LD phase, with $\rho_{3}=\rho_{\text{LD}}$. Further, since the LDW is located inside $T_{2}$, we must have $\rho_{2}=\rho_{2+}$ for $\epsilon_{1}<x<x_{w}$; $\rho_{2}=\rho_{2-}$ for $x_{w}<x<\epsilon_{1}+\epsilon_{2}$. Applying PNC we get
@@
Next, consider an LDW in $T_{3}$. Thus, $\rho_{3}(x)=\rho_{\text{HD}},\;\epsilon_{1}+\epsilon_{2}<x<x_{w}$, $\rho_{3}(x)=\rho_{\text{LD}},\;x_{w}<x<1$. Further, $\rho_{1}(x)=\rho_{\text{HD}}$, $\rho_{2}(x)=\rho_{2+}$, since both $T_{1}$ and $T_{2}$ are in their HD phases. Then applying PNC, we obtain
@@
Our MFT and MCS results on LDWs in $T_{1},\,T_{3}$ are shown in Fig. 8 (top) and Fig. 8 (bottom) respectively. We find reasonable agreement between our MFT and MCS results.
Figure 8: Plot of an LDW in (top) $T_{1}$ in the point defect dominated regime with $p=0.2,n=0.35,q=0.65,L=9000,\epsilon_{1}=\epsilon_{2}=1/3$, and (bottom) $T_{3}$ in the point defect dominated regime with $p=0.2,n=0.65,q=0.65,L=9000,\epsilon_{1}=\epsilon_{2}=1/3$. MFT (blue lines) and MCS (red points) results are shown (see text).
Appendix C LDWs due to the extended defect
@@
$x_{w}\rho_{+}+(\epsilon_{1}-x_{w})\rho_{-}+\frac{\epsilon_{2}}{2}+(1-\epsilon_{1}-\epsilon_{2})\rho_{+}=n.$ (37)
@@
Appendix D Phase diagrams
In this Section, we give the mean-field phase diagrams for our model with unequal segments: We consider (i) $\epsilon_{1}=2/5,\epsilon_{2}=1/5$ and (ii) $\epsilon_{1}=1/4,\epsilon_{2}=1/2$; see Fig. 10 (top) and Fig. 10 (bottom) respectively. These two phase diagrams and the phase diagram in the main text (with $\epsilon_{1}=1/3,\epsilon_{2}=1/3$) have the same topology, with the phase boundaries shifting mildly. Thus the phase diagrams do not sensitively depend on the sizes of the segments, controlled by $\epsilon_{1}$ and $\epsilon_{2}$.
Figure 10: Mean-field phase diagrams with parameters (top) $\epsilon_{1}=2/5,\epsilon_{2}=1/5$ and (bottom) $\epsilon_{1}=1/4,\epsilon_{2}=1/2$. These have the same topology and are very similar to the phase diagram in the main text (with $\epsilon_{1}=1/3,\epsilon_{2}=1/3$)